OpenStudy (anonymous):

IQ scores are normally distributed with a mean of 100 and a standard deviation of 15. If a certain statistician has an IQ of 122, what percent of the population has an IQ less than she does?

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

how od I solve this??

OpenStudy (mathmate):

Are you familiar with the normal distribution table?

OpenStudy (anonymous):

um.. I think so

OpenStudy (mathmate):

such as: http://www.math.unb.ca/~knight/utility/NormTble.htm

OpenStudy (anonymous):

yea.. its so confusing though! anyways what do I have to do?

OpenStudy (mathmate):

normalize the IQ 122 as \( Z=(X-\mu)/\sigma \)

OpenStudy (anonymous):

so.. how/ what would I plug in?

OpenStudy (mathmate):

Look up the table for Z, that will give you the percentage of the population with IQ 122 or less!

OpenStudy (anonymous):

i don't understand? for what number of Z do i look up

OpenStudy (mathmate):

Z=(X−μ)/σ that you calculated with X=122

OpenStudy (anonymous):

ok.. and what abt the denominator symbol??

OpenStudy (mathmate):

In statistics, \( \sigma \) is the standard deviation, \( \mu \) is the mean.

OpenStudy (anonymous):

so Z= (122-100) / 15???

OpenStudy (anonymous):

.9279??? on the z-table? i got 93%?

OpenStudy (mathmate):

Exactly!

OpenStudy (anonymous):

yay that was my guess again!

OpenStudy (anonymous):

i get it now! actually get the z-table! Thank u!

OpenStudy (mathmate):

You can be more accurate by doing a little interpolation between .9279 and .9293

OpenStudy (anonymous):

pls wait. i have to ask u another

OpenStudy (anonymous):

is this an example of data gathering? -finding the mean fasting blood sugar of diabetics

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

hello? are u there

OpenStudy (mathmate):

If "mean fasting blood sugar" is a technical term, then you are gathering some technical information/data. However, it does not seem to be, "fasting blood sugar" is, so you are just working to find the mean of some data before. In this case, this task is not data gathering.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

yeah.. idk what that means either/ i don't THINK its a term...

OpenStudy (anonymous):

fasting blood sugar is

OpenStudy (anonymous):

but not with "mean'